# VL-ICL Bench: The Devil in the Details of Benchmarking Multimodal In-Context Learning

### v1

@@ 4.1 Experiment Setup
Models  We evaluate a diverse family of state-of-the-art models with various sizes (ranging from 7B to 80B) and different LLM backbones on our benchmark. Specifically, for image-to-text VLLMs, we select Open Flamingo (9B) (Awadalla et al., 2023), IDEFICS (9B and 80B) (Laurençon et al., 2023), Otter (9B) (Li et al., 2023b), InterLM-XComposer2 (7B) (Zhang et al., 2023), LLaVA-Next (Vicuna-7B) (Liu et al., 2024a), Qwen-VL-Chat (9B) (Bai et al., 2023), and Emu2-Chat (34B) (Sun et al., 2023a). For text-to-image VLLMs, we use GILL (7B) (Koh et al., 2023), SEED-LLaMA (8B, 14B) (Ge et al., 2024), Emu1 (14B) (Sun et al., 2023b), Emu2-Gen (34B) (Sun et al., 2023a). We also evaluate GPT4V (OpenAI, 2023) on our benchmark. We use officially released model weights or GPT4 API and adopt greedy decoding for reproducibility. All experiments are conducted using three different random seeds and we report the average performance. A100-80GB GPUs are used for experiments.
Prompt  For consistency, we employ the following prompt format for in-context learning. Additionally, we investigate the impact of various prompt formats, with detailed findings presented in the supplementary materials.
[Task Description] Support Set: [Image][Question][Answer] (n-shot) Query: [Image][Question] Prediction: [Answer]
Figure 4: VL-ICL Bench results. Top two rows: Image-to-Text. Bottom: Text-to-Image tasks.
@@ 4.2 Main Results
Figure 5: Illustration of how the best models perform on our benchmark. Left: By dataset separately. Right: By capability evaluated, averaging over datasets.
@@
The main results for VL-ICL Bench are presented in Fig. 4 including a breakdown over shots, and summarised for the 2-shot case as a radar plot over tasks and capabilities in Fig. 5. We make the following observations: (1) VLLMs demonstrate non-trivial in-context learning on VL-ICL Bench tasks. Unlike the common VQA and captioning benchmarks (Fig. 2), our tasks have low zero-shot performance and in every task at least one model shows a clear improvement in performance with number of shots. Thus, ICL capability is now indeed being demonstrated and exploited. (2) VLLMs often struggle to make use of a larger number of ICL examples. For several tasks and models performance increases with the first few shots; but the increase is not monotonic. Performance often decreases again as we move to a larger number of shots (e.g., GPT4V CLEVR Count Induction; InternLM-XComposer2 Operator induction; IDEFICS-80B Interleaved operator induction). Models are eventually confused by this greater number of images and tokens, rather than exploiting them to learn the task at hand. This is exacerbated by the difficulty of extrapolation over context length and number of input images, which for higher-shot ICL becomes greater than the context length and image number used for VLLM training. This shows an important limit of the current state-of-the-art in ICL: Future models must support longer contexts and more images to benefit from larger support sets. (3) GPT4V is the best overall image-to-text model. Among all of the models GPT4V is the strongest all round (but surpassed by some such as OpenFlamingo in low-shot MiniImageNet). (4) Zero-shot performance is not strongly indicative of ICL ability. LLaVA-Next-7B (Liu et al., 2024a) is perhaps worst overall on VL-ICL Bench, which is surprising as it is a state-of-the-art open-source model in mainstream zero-shot benchmarks. This is due to point (2): Its training protocol uses one image at a time, and it uses a large number of tokens per image – thus ICL requires it to extrapolate substantially in input image number and token number, which it fails to do. (5) There is No clear winner among text-to-image models. However, text-to-image models have more consistent shot scaling than image-to-text models. This is due to training with more diverse interleaved datasets that provide multiple input images per instance, and using fewer tokens per image for better scaling.
@@ 4.3 Additional Analysis
We next use VL-ICL Bench to analyse the role of several challenges and factors influencing ICL performance.
@@
Fast Concept Binding  In our open miniImageNet task, we follow (Tsimpoukelli et al., 2021) to require fast-binding of synthetic concept names so as to purely test models’ ICL ability, without confounding by VLLMs’ zero-shot ability to associate visual concepts with names. Fig. 6 compares the fast and real-world miniImageNet recognition, where we see the fast-binding case is more challenging.
Figure 6: Comparison of fast binding vs real-names version of MiniImageNet. Fast-binding has zero accuracy for zero-shot inference, unlike the real-names version. It is thus solely dependent on ICL for success.
@@
Direct Comparison of Multimodal and Text ICL  We can disentangle the role of text versus image inputs for some image-to-text VL-ICL Bench tasks, where we can easily provide a semantically equivalent text input describing the image, in place of image tokens. Fig. 7 shows a comparison between image-input vs text-input for count induction, operator induction, and interleaved operator induction tasks. With text input, performance grows much more sharply and consistently with the number of shots. This is attributable to both (i) reduction of perception difficulty, and (ii) reduction in the total number of tokens compared to image input.
@@
Figure 7: Comparison of multimodal (dashed line) and text (solid line) ICL. Performance increases more sharply and consistently with text inputs.
Scaling with Number of Shots  As discussed in Sec. 4.2, the various models exhibit different scaling abilities with respect to number of shots. We summarise their scaling ability by aggregating over tasks and reporting the average accuracy increment in each shot increment Fig. 8. Evidently, VLLMs vary in their accuracy increment per shot, and how well they can extract knowledge from a growing number of shots.
Figure 8: Aggregate analysis across datasets to find the average performance increments when more shots are added. The individual bars of one color correspond to improvements from 0$\rightarrow$1, 1$\rightarrow$2 and 2$\rightarrow$4.
4.4 Qualitative Analysis
We also include a qualitative analysis, where we analyse the impact of using more support examples on the quality of the output. We analyse text-to-image tasks in Fig. 9, using Emu2-Gen model. For text-to-image MiniImageNet the model should learn from the support examples that the artificial names slation and shously correspond to a lion and a school bus, respectively. Emu2-Gen is able to do it to a certain extent, but may get confused by additional support examples as more support examples are not necessarily helpful. In CoBSAT, the support set induces that the animal should have a glacier and desert background. With no support examples the model only displays the animal, but with more support examples it learns that it should use glacier background in the first example. In the second example, the model is able to capture that it should use desert background, but is less successful in showing the required animal – zebra. The quality of the generated images is not necessarily better with more support examples.
CLEVR Count Induction  In many cases the model rephrases the question, while in others it says that e.g. the described object is present. Such behaviour is more common with fewer or no support examples. With more support examples the model learns to predict a count but gets incorrect answer. It can be because some objects are more difficult to recognize, e.g. if one partially covers another.
Matching MiniImageNet  The models often describe what is shown on one of the images. However, in many cases they simply return the wrong answer, saying no when it should be yes.
@@ 5 Related Work
VLLM Evaluation  With the rapid development of VLLMs, researchers are creating evaluation benchmarks to thoroughly assess the capabilities of VLLMs from diverse perspectives. These evaluations range from zero-shot aggregated benchmarks such as MME (Fu et al., 2023), MMbench (Liu et al., 2023c), and MM-VET (Yu et al., 2023) to datasets designed for fine-tuning on specific aspects, such as visual reasoning (Hudson and Manning, 2019) and knowledge-grounded QA (Lu et al., 2022). They predominantly focus on single-image scenarios, leaving in-context learning evaluation underexplored.
In-Context Learning Evaluation  The term “in-context” has been used in a few ways, including to describe scenarios with interleaved inputs, such as multiple video frames or multi-turn conversations (Li et al., 2023a, b; Zhao et al., 2023; Ge et al., 2024). Although the study of interleaved inputs presents an intriguing subject, it does not align with the core definition of in-context learning that we consider following (Brown et al., 2020; Dong et al., 2022; Min et al., 2022), which involves the emergent ability to learn a function from $x\to y$ from few-shot support input-output pairs. Prior evaluation of ICL (Awadalla et al., 2023; Bai et al., 2023; Laurençon et al., 2023; Sun et al., 2023a) in this sense is limited, and comes with serious drawbacks as discussed in Sec. 2.2. Concurrent to our work, CobSAT (Zeng et al., 2024) introduces a benchmark designed to evaluate in-context learning in text-to-image models, focusing particularly on latent variable induction capabilities. Our work expands upon this by encompassing tasks for both image-to-text and text-to-image generation, assessing a wider array of capabilities (Tab. 1). Additionally, we have incorporated CobSAT as a subset of our benchmark.
@@
6 Conclusion
We have introduced the first comprehensive benchmark suite VL-ICL Bench for multimodal vision-and-language in-context learning with VLLMs. This benchmark suite avoids the issue with the existing mainstream but limited approach to evaluating image-to-text ICL – that ICL provides limited demonstrable benefit over zero-shot inference, and VLLMs learn answer formatting at best rather than any true multimodal capability. In contrast, VL-ICL Bench tests a wide variety of multimodal capabilities including both text-to-image and image-to-text generation, fine-grained perception, rule-induction, reasoning, image interleaving, fast concept binding, long context, and shot scaling. We hope this benchmark will inspire model developers to consider all these capabilities in VLLM development, and inform practitioners about the evolution of what VLLM ICL can and cannot do as the field develops. We also aim to expand our benchmark to incorporate more tasks and models in the future.

### v2

@@ 4.1 Experiment Setup
Models  We evaluate a diverse family of state-of-the-art models with various sizes (ranging from 0.5B to 80B) and different LLM backbones on our benchmark. Specifically, for image-to-text VLLMs, we select Open Flamingo (9B) (Awadalla et al., 2023), IDEFICS (9/80B) (Laurençon et al., 2023), IDEFICS-v2 (8B) (Laurençon et al., 2024), Otter (9B) (Li et al., 2023b), InternLM-XComposer2 (7B) (Zhang et al., 2023), LLaVA-Next (Vicuna-7B) (Liu et al., 2024a), Qwen-VL-Chat (9B) (Bai et al., 2023), Emu2-Chat (34B) (Sun et al., 2023a), VILA (7B) (Lin et al., 2024), Mantis (-Idefics2) (Jiang et al., 2024a), Phi-3-Vision (4B) (Abdin et al., 2024), LongVA (Zhang et al., 2024b), InternLM-XComposer2.5 (7B) (Zhang et al., 2024a), and LLaVA-OneVision (0.5/7/72B) (Li et al., 2024). For Text-to-image VLLMs, we use GILL (7B) (Koh et al., 2023), SEED-LLaMA (8B, 14B) (Ge et al., 2024), Emu1 (14B) (Sun et al., 2023b), Emu2-Gen (34B) (Sun et al., 2023a). We also evaluate GPT4V (OpenAI, 2023) on our benchmark. We use officially released model weights or GPT4 API and adopt greedy decoding for reproducibility. All experiments are conducted using three different random seeds and we report the average performance.
Evaluation Metrics  All our experiments evaluate test accuracy as a function of the number of shots (cf: Fig. 3 and Fig. 5). To summarise these curves for easy comparison we use three main metrics: zero-shot accuracy, peak (max.) accuracy over all shots and ICL efficiency (the area under the accuracy vs shots curve above the zero-shot starting point, normalized over the whole area). For text-to-image models, we employ LLaVA-Next-7B as the judge model to determine whether the generated images are correct (e.g., whether contain the target attribute). Implementation details are detailed in Appendix A.
Prompt  For consistency, we employ the following standard prompt format for in-context learning.
{mdframed}
[backgroundcolor=gray!20] [Task Description]
Support Set: [Image][Question][Answer] (n-shot)
Query: [Image][Question]
Prediction: [Answer]
@@ 4.2 Main Results
The main results for VL-ICL Bench are illustrated for selective subsets in Fig. 5 and summarised quantitatively in Tab. 2 and Tab. 3 for I2T and T2I benchmarks respectively. We make the following observations: (1) VLLMs demonstrate non-trivial in-context learning on VL-ICL Bench tasks. Unlike the common VQA and captioning benchmarks (Fig. 3), our tasks have low zero-shot performance and in every task at least one model shows a clear improvement in performance with number of shots. Thus, ICL capability is now indeed being demonstrated and exploited. (2) VLLMs often struggle to make use of a larger number of ICL examples. For several tasks and models performance increases with the first few shots; but the increase is not monotonic. Performance often decreases again as we move to a larger number of shots (e.g., GPT4V CLEVR Count Induction; InternLM-XComposer2 Operator induction in Fig. 5). More extremely, some models obtain negative impact from more shots, leading to negative ICL efficiency in Tab. 2. We attribute this to difficulty of dealing with the larger number of images and tokens confusing the model and overwhelming the value of additional training data. It is exacerbated by the difficulty of extrapolation over context length and number of input images, which for higher-shot ICL becomes greater than the context length and image number used for VLLM training. This shows an important limit of the current state-of-the-art in ICL: Future models must support longer contexts and more images to benefit from larger support sets. (3) LLaVA-OneVision 72B is the best overall image-to-text model, closely followed by GPT4V. (4) Zero-Shot Performance is not strongly indicative of ICL ability. For example LLaVA-Next-7B (Liu et al., 2024a) is perhaps one of the worst overall on VL-ICL Bench, which is surprising as it is a strong model in mainstream zero-shot benchmarks. This is due to point (2): Its training protocol uses one image at a time, and it uses a large number of tokens per image – thus ICL requires it to extrapolate substantially in input image number and token number, which it fails to do. (5) There is No clear winner among text-to-image models. However, text-to-image models have more consistent shot scaling than image-to-text models. This is due to training with more diverse interleaved datasets that provide multiple input images per instance, and using fewer tokens per image for better scaling.
We remark that our zero-shot and ICL efficiency summary metrics (Tabs 2,  3) can sum to one at most. These metrics are visualised together in Fig. 14, along with the best possible Pareto-front. The differing degree to which each model/dataset relies on ZS vs ICL performance is easily visible.
Table 2: Average zero-shot, peak and efficiency scores (%, $\uparrow$) of different models on VL-ICL Bench image-to-text datasets. LLaVA-OneVision-72B and GPT4V show the best ICL abilities.
Model Fast Open-Ended MiniImageNet CLEVR Count Induction Operator Induction TextOCR Interleaved Operator Induction Fast Matching MiniImageNet Avg. Rank
Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff.
OpenFlamingo-9B 00.0 58.2 -46.2 00.0 18.8 -16.6 05.0 07.8 0-1.1 00.0 00.0 -00.0 00.0 08.9 -04.7 00.0 29.1 -20.2 09.3 13.3 08.7
IDEFICS-9B 00.0 59.2 -42.1 00.0 30.3 -26.5 11.7 14.4 0-1.5 16.5 28.0 -06.6 15.0 15.0 0-8.7 00.0 00.1 -00.0 07.7 10.8 08.7
IDEFICS-80B 00.0 62.5 -42.5 00.0 32.4 -29.6 13.3 21.7 -04.3 20.0 29.5 -06.1 25.0 36.7 -02.7 00.0 28.3 -22.3 07.0 07.8 05.8
IDEFICS2-8B 00.0 61.5 -37.3 02.0 51.5 -43.0 36.7 47.2 -04.6 23.0 30.3 -03.5 38.3 38.3 -18.1 00.0 25.2 -10.3 03.0 05.5 08.5
Otter 00.0 28.5 -20.6 00.0 08.3 -05.3 21.7 21.7 -10.0 00.0 00.8 -00.5 08.3 09.4 0-1.0 00.0 00.0 -00.0 08.0 14.5 13.0
InternLM-X2 00.0 50.3 -34.1 01.8 26.0 -19.4 26.1 40.0 -10.0 08.7 16.0 -03.3 28.3 28.3 -18.2 00.0 49.9 -25.3 05.5 09.3 10.0
Qwen-VL-Chat 00.0 58.0 -37.2 00.0 30.2 -26.1 15.0 25.0 -03.8 04.8 24.2 -16.1 16.7 16.7 0-8.2 00.0 00.3 -00.1 07.7 10.8 08.3
LLaVA-Next-7B 00.0 37.2 -29.4 00.0 25.2 -14.5 10.6 10.6 0-6.8 24.7 24.7 -23.0 13.9 13.9 0-7.8 00.0 00.0 -00.0 07.3 13.7 13.2
Emu2-Chat 00.0 29.3 -21.6 05.3 17.7 -09.1 28.6 28.6 0-6.7 25.8 36.5 -05.7 26.7 26.7 -13.2 00.0 40.2 -32.9 04.0 10.2 11.0
VILA-7B 00.0 38.2 -32.3 03.5 34.3 -27.5 28.3 28.3 -18.9 28.0 30.2 0-3.7 28.3 28.3 -15.7 00.0 49.9 -44.6 03.7 08.0 10.8
Mantis-Idefics2 00.0 84.3 -55.2 31.4 40.2 0-3.9 16.7 16.7 0-2.9 21.0 27.8 -04.1 30.0 30.0 -22.3 00.0 51.1 -43.8 04.5 06.7 09.7
Phi3-Vision 00.0 50.0 -33.3 00.0 34.5 -20.4 00.0 54.4 -46.8 08.0 41.5 -28.9 00.0 27.8 -18.9 00.0 50.8 -42.4 09.2 06.3 05.3
LongVA-7B 00.0 52.7 -39.9 00.0 29.2 -21.4 18.3 18.3 -11.9 26.0 26.0 -11.0 10.0 10.0 0-5.3 00.0 50.0 -45.0 06.3 10.5 09.3
LLaVA-OneVision-72B 00.0 98.7 -68.4 00.5 42.3 -32.7 33.3 75.6 -29.8 48.5 51.7 -02.4 38.3 47.8 -02.8 00.0 70.3 -50.8 02.5 01.7 03.8
InternLM-X2d5 00.0 91.0 -62.7 06.0 63.2 -47.5 36.7 41.7 -00.3 32.0 42.5 -04.2 38.3 38.3 0-8.0 00.0 46.8 -26.1 01.5 03.7 05.8
GPT4V 00.0 78.0 -41.8 06.0 42.0 -29.0 24.0 92.0 -59.0 39.3 50.0 -07.2 36.0 74.0 -32.2 00.0 58.2 -38.3 02.8 02.3 03.8
Table 3: Average zero-shot, peak and efficiency scores (%, $\uparrow$) of different models on VL-ICL Bench text-to-image datasets.
Model Text-to-Image Fast MiniImageNet CoBSAT Fast Counting Fast Attribute Matching Avg. Rank
Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff.
GILL 00.0 16.0 -13.6 02.7 12.3 -07.1 34.2 34.2 -26.3 20.5 20.5 -13.2 03.5 04.8 04.8
SEED-LLaMA-8B 00.0 16.5 -13.3 00.5 33.7 -24.5 34.2 56.7 -12.7 21.0 34.3 -10.5 03.5 02.5 02.2
SEED-LLaMA-14B 00.8 21.2 -16.3 05.5 43.8 -30.7 41.3 51.6 -01.9 22.4 35.6 -08.4 02.0 02.0 02.5
Emu1-Gen 00.5 31.5 -22.5 00.3 09.7 -07.1 33.2 47.8 -08.2 23.5 29.1 -03.1 03.5 03.8 02.8
Emu2-Gen 00.0 37.0 -28.6 08.7 28.7 -15.6 44.2 59.2 -05.2 26.5 34.0 0-0.2 01.5 02.0 02.8
Figure 5: Illustrative VL-ICL Bench results. Our tasks better exploit and evaluate ICL compared to the mainstream in Fig. 3. They have lower zero-shot accuracy and more substantial gains with shots.
@@ 4.3 Additional Analysis
We next use VL-ICL Bench to analyse several challenges and factors influencing ICL performance.
@@
Fast Concept Binding  In our open miniImageNet task, we follow Tsimpoukelli et al. (2021) to require fast-binding of synthetic concept names so as to purely test models’ ICL ability, without confounding by VLLMs’ zero-shot ability to associate visual concepts with names. Tab. 4 compares the fast and real-world miniImageNet recognition. Our fast-binding case is much more challenging.
Table 4: Average zero-shot, peak and efficiency scores (%, $\uparrow$) on fast open-ended and real-world MiniImageNet dataset. Real-world version is much easier, with larger zero-shot and peak accuracies.
Dataset IDEFICS-80B Otter InternLM-X2 Qwen-VL-Chat LLaVA-Next-7B GPT4V
Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff.
Fast Open-Ended 00.00 62.50 42.54 00.00 28.50 20.62 00.00 50.33 34.10 00.00 58.00 37.22 00.00 33.67 14.57 00.00 78.00 41.80
Real-World 30.50 94.67 43.05 13.00 61.00 38.75 20.00 67.00 35.12 32.17 88.33 30.81 20.50 64.50 13.43 48.00 90.00 27.00
@@
Direct comparison of multimodal and text ICL  We can disentangle the role of text versus image inputs for some image-to-text VL-ICL Bench tasks, where we can easily provide a semantically equivalent text input describing the image, in place of image tokens. Fig. 6 shows a comparison between image-input vs text-input for count induction, operator induction, and interleaved operator induction tasks. With text-input, performance grows much more sharply and consistently with number of shots. This is attributable to both (i) reduction of perception difficulty, and (ii) reduction in the total number of tokens compared to image input.
@@
Figure 6: Comparison of multimodal (dashed line) and text (solid line). Performance increases more sharply and consistently with text inputs, highlighting the greater difficulty of multimodal ICL.
Disentangling context length and in-context learningOne direct factor limiting VL ICL is the model’s context length, as images translate to a large number of tokens (e.g., in LLaVA, one image translates to 576 tokens, causing an 8-shot setting to exceed a 4k context window). This makes it difficult to disentangle whether the limitation on VL-ICL performance arises from context length constraints or inherent challenges in the models’ ability to perform ICL. To further understand this issue, we apply SelfExtend (Jin et al., 2024), a training-free position encoding extrapolation strategy that has been shown to effectively extend the context length of LLMs. We use SelfExtend to extend the context length of LLaVA and VILA from 4k to 16k. As shown in Tab. 5 (Fig. 15 in the Appendix), though it can help in some cases, performance often remains similar, suggesting that long-context capabilities are necessary but not sufficient for improving ICL. Furthermore, even recent models with 128k context windows, like LLaVA-OneVision, do not always show improved performance with more shots as seen in Tab. 2. This highlights that VL-ICL is a more complex challenge, requiring an understanding of interleaved few-shot examples, which cannot be solved solely by increasing context length.
Table 5: Comparison of models with and without context extension strategy (SelfExtend). While it is helpful in some cases, context extension does not necessarily improve the performance of ICL.
Dataset Fast Open-Ended MiniImageNet CLEVR Count Induction Operator Induction TextOCR
Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff. Z.s. Pk. Eff.
LLaVA-Next-7B (w/o SelfExtend) 00.0 37.2 -29.4 00.0 25.2 -19.3 10.6 10.6 0-6.8 24.7 24.7 -23.0
LLaVA-Next-7B (w/ SelfExtend) 00.0 51.0 -38.9 00.0 29.0 -25.4 11.7 11.7 0-5.8 26.0 26.0 -23.7
VILA-7B (w/o SelfExtend) 00.0 38.2 -32.3 03.5 34.3 -27.5 28.3 28.3 -18.9 28.0 30.2 0-3.7
VILA-7B (w/ SelfExtend) 00.0 54.0 -40.0 04.0 34.8 -27.5 28.3 28.3 -20.3 28.0 29.7 0-4.5
Emergent threshold of multimodal ICLWe further investigate whether there is a threshold in model size at which emergent ICL abilities appear, as described by Wei et al. (2022a) for LLMs. To test this, we use the LLaVA-OneVision series (Li et al., 2024), which share the same architecture and training data but differ in LLM size (0.5B, 7B, and 72B parameters). Notably, we find that the 72B model demonstrates emergent ICL abilities. As shown in Figure 7, the 72B model improves with more shots across all tasks, particularly in Interleaved Operator Induction and Fast Matching MiniImageNet, where the 0.5B and 7B models perform worse than chance as shots increase. This indicates that the 72B model understands the tasks, while the smaller models fail, highlighting the impact of model size on ICL and the presence of an emergent threshold.
Figure 7: Comparison of different model sizes. Dashed line indicates random chance performance assuming available options (the models do open-ended generation so can be worse than random).
Additionally, we provide thorough analyses in the Appendix B and C on various factors, including scaling to many shots (up to 64), ICL efficiency versus zero-shot performance, chain-of-thought prompting, repeating the support set, varying levels of task descriptions, and qualitative results.
@@ 5 Related Work
VLLM Evaluation  With the rapid development of VLLMs, researchers are creating evaluation benchmarks to thoroughly assess their capabilities from diverse perspectives. These evaluations range from zero-shot aggregated benchmarks like MME (Fu et al., 2023), MMbench (Liu et al., 2023c), and MM-VET (Yu et al., 2023) to datasets designed for fine-tuning on specific aspects, such as visual reasoning (Hudson and Manning, 2019) and knowledge-grounded QA (Lu et al., 2022). They predominantly focus on single-image, leaving in-context learning evaluation underexplored.
In-Context Learning Evaluation  The term “in-context” has been used in a few ways, including to describe scenarios with interleaved inputs, such as multiple video frames or multi-turn conversations (Li et al., 2023a, b; Zhao et al., 2023; Ge et al., 2024). Although the study of interleaved inputs presents an intriguing subject, it does not align with the core definition of in-context learning that we consider following (Brown et al., 2020; Dong et al., 2022; Min et al., 2022), which involves the emergent ability to learn a function from $x\to y$ from few-shot support input-output pairs. Prior evaluation of ICL (Awadalla et al., 2023; Laurençon et al., 2023; Sun et al., 2023a; Baldassini et al., 2024; Chen et al., 2023) in this sense is limited, and comes with serious drawbacks as discussed in Sec. 2.2. Concurrent to our work, CobSAT (Zeng et al., 2024) introduces a benchmark designed to evaluate in-context learning in text-to-image models, focusing particularly on latent variable induction capabilities. Our work spans a much wider range of tasks and capabilities, encompasses both image-to-text and text-to-image generation and subsumes CobSAT as one of our ten total benchmarks (Tab. 1). The concurrent work of Jiang et al. (2024b) studies ICL shot-scaling, but only for I2T recognition tasks - in contrast to our wide range of task types (Tab. 1), and without elucidating the limitations of mainstream VQA/captioning paradigm as we have in Sec. 2.2.
@@
6 Discussion
We have introduced the first comprehensive benchmark suite VL-ICL Bench for multi-modal vision-and-language in-context learning. This benchmark suite avoids the issue with the existing mainstream but limited approach to evaluating image-to-text ICL – that ICL provides limited demonstrable benefit over zero-shot inference, and VLLMs learn answer formatting at best rather than any true multi-modal capability. In contrast, VL-ICL Bench tests a wide variety of multi-modal capabilities including both text-to-image and image-to-text generation, fine-grained perception, rule-induction, reasoning, image interleaving, fast concept binding, long context, and shot scaling. We hope this benchmark will inspire model developers to consider all these capabilities in VLLM development, and inform practitioners about the evolution of what VLLM ICL can and cannot do as the field develops. One limitation of this work is that we evaluate only a small number of text-to-image models, in contrast to the more comprehensive set of image-to-text models. This is due to the limited availability of VLLMs capable of handling interleaved inputs and generating images. Developing such models presents a promising direction for future research.